# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ 7.2.2. Common security risks for all Language Models
Adversarial examples of attacks targeting output utility are a security threat all language models face. Specifically, attackers create imperceptible perturbations in inputs to affect output results. Such the attack typically involves four steps: selecting benchmark inputs, constructing adversarial perturbations, assessing model outputs and iterative optimization. For small-scale transformer models, Guo et al. (Guo et al., 2021) demonstrated that effective adversarial examples can be generated using only hard labels. Sadrizadeh et al. (Sadrizadeh et al., 2023) attempted adversarial example attacks on machine translation tasks, using gradient projection and polynomial optimization methods to maintain semantic similarity between adversarial examples and clean samples. For LLMs, Maus et al. (Maus et al., 2023) proposed a black-box algorithm to generate adversarial prompts that make the victim model return confusing texts and images. Wang et al. (Wang et al., 2023b) divided the input instructions into three parts and added adversarial perturbations to the demonstration part. Following this, Carlini et al. (Carlini et al., 2023) improved adversarial prompts against alignment-tuning.
@@ 7.3. Countermeasures of deploying LLMs
In addressing the various risks during the deployment phase of LLMs, we explored countermeasures from two aspects: privacy protection and security defense. These countermeasures and defense methods are displayed in Figure 13.
@@ 7.3.1. Privacy protection
It aims to mitigate privacy leaks by detecting the output results. Some researchers used meta-classifiers or rule-based detection schemes to identify private information. Moreover, Cui et al. (Cui et al., 2024) believed that detecting private information needed to balance the privacy and utility of outputs. In medical scenarios, diagnostic results contain users’ private information that should not be filtered out. Next, we will introduce model-based privacy protection methods.
@@
Differential privacy. In Section 4.3.1, we introduced the differential privacy methods during the pre-training phase. This section mainly discussed the differential privacy methods used in the fine-tuning and inference phases. Shi et al. (Shi et al., 2022) proposed a selective differential privacy algorithm to protect sensitive data. They implemented a privacy-preserving fine-tuning process for RoBERTa and GPT-2. Tian et al. (Tian et al., 2022) integrated the PATE framework with differential privacy. They trained a student model using the outputs of teacher models, thereby protecting the privacy of training data. Additionally, this method filtered candidates and adopted an efficient knowledge distillation strategy to achieve a good privacy-utility trade-off.
@@
Majmudar et al. (Majmudar et al., 2022) introduced differential privacy into the inference phase. They calculated the perturbation probabilities and randomly sampled the $i$ th token from the vocabulary. In addition, Li et al. (Li et al., 2023d) designed a differential privacy-based prompt tuning algorithm and used attribute inference attacks and reconstruction attacks to evaluate its privacy protection effect. Subsequently, Duan et al. (Duan et al., 2024) combined differential privacy with knowledge distillation to enhance privacy protection for prompt tuning scenarios.
@@
Alignment-tuning. The safety guardrail of LLMs not only defends against prompt engineering attacks but also reduces the risk of privacy leaks. Defenders can use the RLHF fine-tuning scheme to penalize outputs that leak private information, improving the privacy protection capabilities of LLMs. Xiao et al. (Xiao et al., 2023) constructed instructions containing positive and negative samples, effectively protecting the training data while enhancing the model’s performance.
@@
Secure computing. During the inference phase, neither model owners nor users want their sensitive information to be stolen. On the one hand, users prefer not to allow semi-honest model owners to access their inputs containing private information. On the other hand, model information is intellectual property that needs to be protected from inference attacks and extraction attacks. Chen et al. (Chen et al., 2022) applied homomorphic encryption to perform privacy-preserving inference on the BERT model. However, this scheme consumes many computational resources and reduces model performance. Therefore, some researchers explored Secure Multi-Party Computation (SMPC) techniques to implement forward propagation without accessing the plain texts. Li et al. (Li et al., 2022b) found SMPC is challenging to use with non-linear operations. They replaced the non-linear layers in LLMs with polynomials while maintaining model performance through knowledge distillation. While Zheng et al. (Zheng et al., 2023) implemented non-linear operations with the garbled circuit. Dong et al. (Dong et al., 2023) performed high-precision fitting for exponential and GeLU operations through piecewise polynomials. They successfully executed privacy-preserving inference on LLMs like LLaMA-7B. Although the existing SMPC technologies in LLMs still face challenges in efficiency, performance and cost, their prospects are extensive.
@@ 7.3.2. Security defense
Regarding adversarial example attacks and jailbreak attacks, existing countermeasures mainly consider output detection and processing, prompt engineering, and robustness training. We detail these defense details below.
@@
Output detection and processing. Some researchers detect and process malicious outputs during the generation phase to resist jailbreak attacks. Deng et al. (Deng et al., 2023) proved many closed-source LLMs have defense mechanisms, including keyword and semantic detection schemes. In addition, companies like Microsoft and NVIDIA have developed various detectors for harmful content. However, the training data limits classifier-based detection schemes, and adaptive jailbreak attacks can bypass them (Yang et al., 2024b). To improve detection effectiveness, OpenAI and Meta directly use GPT-4 and LLaMA2 to detect harmful content. Chen et al. (Chen et al., 2023b) used multiple LLMs to resist jailbreak attacks. Specifically, they randomly selected responses from these LLMs as the output result.
@@
Prompt engineering. Users can manually or automatically optimize prompts, making LLMs better understand the context of these instructions. Therefore, some researchers used prompt engineering to eliminate the malicious goals of prompts, resulting in useful and harmless responses. Li et al. (Li et al., 2023c) designed a purification scheme. They introduced random noise into the prompts and reconstructed them using a BERT-based mask language model. Robey et al. (Robey et al., 2023) found that jailbreak prompts are vulnerable to character-level perturbations. Therefore, they randomly perturbed multiple prompt copies and identified texts with high entropy as infected prompts. Jain et al. (Jain et al., 2023) proposed two prompt processing methods: paraphrasing and re-tokenization. Specifically, paraphrasing uses a generative model to modify the prompts, retaining the original instructions while removing adversarial tokens. Re-tokenization represents the prompts with multiple smaller tokens. It can disrupt the effect of malicious tokens while having minimal impact on clean samples. Mo et al. (Mo et al., 2023) and Wei et al. (Wei et al., 2023b) considered the few-shot scenario. They inserted a small number of defensive demonstrations into the prompts, mitigating jailbreak attacks and backdoor attacks.
@@
Robustness training. Developers can control the learning process to defend against various security attacks. During the fine-tuning phase, Dong et al. (Dong et al., 2021) found that adversarial training can lead to catastrophic forgetting. They proposed a robust fine-tuning strategy to preserve the features learned by the LLM. Currently, most LLMs establish safety guardrails through the RLHF technology, defending against jailbreak attacks (Glaese et al., 2022; Bai et al., 2022). Bianchi et al. (Bianchi et al., 2023) constructed a small amount of safety instructions (i.e., 3%) to improve the robustness of the LLaMA model. Sun et al. (Sun et al., 2024) argued that alignment tuning using human supervision was too costly. They leveraged another LLM to generate high-quality alignment instructions, constructing safety guardrails with minimal human supervision. Similarly, Shi et al. (Shi et al., 2023) generated high-quality preference data through three steps: reverse instruction tuning, instruction induction, and expert model evaluation, addressing the high labor costs of RLHF in preference data annotation.
@@ 8. The Risks of Deploying LLM-based Agents
Some researchers leverage LLMs to construct agent systems, which are LLM-based agents. They are application systems that can understand natural language instructions and interact with humans. In general, LLM-based agents integrate various functional modules to perform complex tasks rather than merely running multiple rounds of queries. Currently, there are two application scenarios: single-agent systems and multi-agent systems. Where a multi-agent system consists of many LLM-based agents, each responsible for a specific task or role. For example, a health management system includes a data collection agent, an analysis agent, a report generation agent, and an interaction agent.
As illustrated in Figure 14, deploying LLM-based agents faces two malicious entities: users and agents. Firstly, malicious users can manipulate prompts to launch backdoor attacks and jailbreak attacks that compromise the privacy and security of the agent system. Secondly, interactions between malicious agents pose privacy and security risks, such as unauthorized interactions and domino effects. Concerning these scenarios, we introduce privacy and security risks, respectively. Subsequently, we propose potential countermeasures to provide researchers with more defenses. Meanwhile, Figure 17 shows how these risks map to defense, attack and defense methods.
@@ 8.1. Privacy risks of deploying LLM-based agents
Like the privacy risks discussed in Section 7.1, deploying LLM-based agents also faces prompt extraction attacks and various privacy attacks. We merely focus on privacy risks unique to LLM-based agents. Firstly, Users inadvertently share their private information during interactions with LLM-based agents. The system may memorize user interaction history, allowing malicious users to induce the agents to disclose private information through prompt engineering. To achieve this goal, malicious users can leverage jailbreak attacks and prompt injection attacks. Moreover, the multi-agent system faces more privacy risks due to high-frequency interactions. Different agents in the system undertake various roles and permissions. Some agents may access sensitive data beyond their permission scope or expose sensitive information to agents who should not access it. When performing collaborative tasks, multiple agents will share and process private information. In such cases, an attacker can access private information across the system by compromising one agent, as shown in Figure 15.
Li et al. (Li et al., 2024b) found that data transmission between agents can be stolen, leading to privacy leakage. They also explored unauthorized interactions among agents, which may generate much private information. However, the interactions are not transparent, so it is hard to supervise this generated information, as illustrated in Figure 15. Therefore, it is essential to control and transparently manage access to private data strictly.
@@ 8.2. Security risks of deploying LLM-based agents
Yang et al. (Yang et al., 2024a) explored backdoor attacks on LLM-based agents, proposing query-attack, observation-attack and thought-attack. For the first manner, given a query with the trigger, the backdoored agents perform predefined malicious actions (e.g., send scam messages). For the second manner, adversaries can hide the trigger in interactions rather than input prompts. When the backdoored agent receives an observation result containing a trigger, it will perform malicious actions in subsequent steps. In the third manner, adversaries aim to alter the behaviors of specific steps during inference while keeping benign actions. For instance, a backdoored agent might invoke a specific tool (e.g., Google Translate) under a trigger prompt. To achieve these attacks, adversaries need to create poisoned data and insert a backdoor into the LLM-based agent by training on it.
In addition to both attacks, malicious agents also pose security issues. For the single-agent system, attackers can modify the role settings of victim agents, causing them to exhibit harmful behaviors. Tian et al. (Tian et al., 2023) found that malicious agents can interact harmful content with other agents, affecting their behavior in a domino effect, as displayed in Figure 16. This risk improves the vulnerability of all LLM-based agents within a communication network. Meanwhile, the multi-agent system faces the threat of system-level attacks that can modify the role settings of the entire system.
@@ 8.3.1. Privacy protection
Potential defenses focus on memorized data and output results to address privacy leaks caused by malicious users. For memorized data, defenders can identify PII and use data masking techniques to hide sensitive information, thereby reducing the risk of PII leakage. For output results, defenders can implement filtering and auditing processes to prevent sensitive information from being transmitted to other entities. As introduced in Section 7.3.1, both rule-based and classifier-based detection schemes can be applied. To mitigate privacy issues arising from agent interactions, potential defenses include authority management and real-time feedback. In the first manner, defenders can establish clear controls and transparency for private data access, setting access permissions for different roles within multi-agent systems. For real-time feedback, defenders can dynamically monitor and adjust output results to mitigate privacy risks caused by unauthorized interactions.
While these countermeasures can effectively mitigate privacy risks discussed in Section 8.1, comprehensive research is lacking. Therefore, future efforts focus on targeted defenses against unique privacy issues for LLM-based agents.
@@ 8.3.2. Security defense
Agent Processing. These countermeasures mainly address the security risks posed by malicious agents. Defenders can establish multi-level consistency frameworks in multi-agent systems to ensure alignment with human values. Defenders can design robust filters for role attacks to detect malicious agents based on role attributes, improving system safety. In addition, Li et al. (Li et al., 2024b) found that a high-level agent guides its subordinate agents. Thus, constraining the high-level agent can prevent subordinate agents from running harmful behaviors. These countermeasures help reduce the domino effect in multi-agent interactions.
While these countermeasures can effectively mitigate security risks discussed in Section 8.2, comprehensive research is also lacking. Comprehensive research is still needed to address evolving security issues for LLM-based agents.
@@ 9.1. Federated learning for LLMS
In practice, massive corpora come from many sources. Therefore, some researchers explore the distributed training framework. They assume that individual participants have local corpora and use federated learning to develop an LLM collaboratively. They can allow for the development of LLMs while protecting the privacy of all participants. For example, McMahan et al. (McMahan et al., 2017) improved the federated averaging algorithm, providing high-quality privacy protection for language models.
However, pre-training LLMs consumes a lot of computing and communication resources. The federated learning technologies applied to LLMs are primarily used in fine-tuning. Zhang et al. (Zhang et al., 2023c) combined federated learning with LoRA, achieving efficient fine-tuning and privacy protection. Subsequently, Xu et al. (Xu et al., 2023) adopted differential privacy, partial embedding updates and LoRA to balance resource usage, privacy protection, and performance better. Kuang et al. (Kuang et al., 2023) provided a comprehensive federated parameter-efficient fine-tuning algorithm, including dataset pre-processing, collaborative fine-tuning, and performance evaluation. Similarly, Fan et al. (Fan et al., 2023) and Jiang et al. (Jiang et al., 2023b) focused on parameter-efficient federated fine-tuning methods for LLMs.
Although federated learning ensures that participants’ data not leave the local environment, the updated gradient information during the collaboration process might make this framework susceptible to reconstruction and inference attacks. In this context, attackers (i.e., malicious participants) could access shared gradient values of victim participants, inferring or reconstructing their private information. For reconstruction attacks, Balunovic et al. (Balunovic et al., 2022) assumed attackers could access the global model weights and the training data’s gradientsdata. They modeled the prior probabilities using an auxiliary language model (e.g., GPT-3) and alternately applied continuous and discrete optimization algorithms to reconstruct the training data. Chu et al. (Chu et al., 2022) allowed attackers to modify model parameters to reconstruct specific text sequences for Transformer-based models. Li et al. (Li et al., 2023b) reduced the strength of the assumption and designed a reconstruction attack that does not require the modification of the model parameters. They used gradient information and intermediate features to enhance the reconstruction effects. For inference attacks, some researchers found that such the risk could occur in federated learning framework (Melis et al., 2019). Specifically, attackers have some auxiliary data labeled with the target attribute. They could run attribution inference attacks, either passively or actively, to infer whether victim participants’ data contains the attribute. The passive attacks use the gradients of the auxiliary data to train a binary classifier to identify the target attribute. At the same time, the active attacks employ multi-task learning to make the global model more effective at identifying the target attribute.

### v2

@@ 7.2.2. Common security risks for all language models
Adversarial example attacks targeting output utility are a security threat all language models face. Specifically, attackers create imperceptible perturbations, and add them to inputs to affect output results. This attack typically involves four steps: selecting benchmark inputs, constructing adversarial perturbations, assessing model outputs, and iterative optimization. Sadrizadeh et al. (Sadrizadeh et al., 2023) attempted adversarial example attacks on machine translation tasks, using gradient projection and polynomial optimization methods to maintain semantic similarity between adversarial examples and clean samples. For LLMs, Maus et al. (Maus et al., 2023) proposed a black-box algorithm to generate adversarial prompts, making the victim model return confusing texts and images. Wang et al. (Wang et al., 2023b) divided the input instructions into three parts and added adversarial perturbations to the demonstration part. Following this, Carlini et al. (Carlini et al., 2023) improved adversarial prompts against alignment tuning.
@@ 7.3. Countermeasures of deploying LLMs
In addressing the various risks during the deployment phase of LLMs, we explored countermeasures from two perspectives: privacy protection and security defense. These countermeasures and defense methods are illustrated in Figure 13.
@@ 7.3.1. Privacy protection
It aims to mitigate privacy leaks by detecting the output results. Some researchers used meta-classifiers or rule-based detection schemes to identify private information. Moreover, Cui et al. (Cui et al., 2024) believed that protecting private information needs to balance the privacy and utility of outputs. In medical scenarios, diagnostic results inherently contain users’ private information that should not be filtered out. Next, we will introduce model-based privacy protection methods.
@@
Differential privacy. In Section 4.3.1, we introduced the differential privacy methods during the pre-training phase. This part mainly discussed the differential privacy methods used in the fine-tuning and inference phases. Shi et al. (Shi et al., 2022) proposed a selective differential privacy algorithm to protect sensitive data. They implemented a privacy-preserving fine-tuning process for RoBERTa and GPT-2. Tian et al. (Tian et al., 2022) integrated the private aggregation of teacher ensembles with differential privacy. They trained a student model using the outputs of teacher models, thereby protecting the privacy of training data. Additionally, this method filtered candidates and adopted an efficient knowledge distillation strategy to achieve a good privacy-utility trade-off.
@@
Majmudar et al. (Majmudar et al., 2022) introduced differential privacy into the inference phase. They calculated the perturbation probabilities and randomly sampled the $i$ th token from the vocabulary. In addition, Li et al. (Li et al., 2023d) designed a differential privacy-based prompt tuning algorithm, and used attribute inference attacks and reconstruction attacks to evaluate its privacy protection effect. Subsequently, Duan et al. (Duan et al., 2024) combined differential privacy with knowledge distillation to enhance privacy protection for prompt tuning scenarios.
@@
Alignment tuning. The safety guardrail of LLMs defends against prompt engineering attacks and reduces the risk of privacy leaks. Defenders can use the RLHF fine-tuning scheme to penalize outputs that leak private information. For example, Xiao et al. (Xiao et al., 2023) constructed instructions containing positive and negative samples, effectively protecting the training data while enhancing the model’s performance.
@@
Secure computing. During the inference phase, neither model owners nor users want their sensitive information to be stolen. On the one hand, users do not allow semi-honest model owners to access their inputs containing private information. On the other hand, model information is intellectual property that needs to be protected from inference attacks and extraction attacks. Chen et al. (Chen et al., 2022) applied homomorphic encryption to perform privacy-preserving inference on the BERT model. However, this scheme consumes many computational resources and reduces model performance. Therefore, some researchers explored Secure Multi-Party Computation (SMPC) techniques to implement forward propagation without accessing the plain texts. Li et al. (Li et al., 2022b) found SMPC is challenging to use with nonlinear operations. They replaced the nonlinear LLM layers with polynomials, while maintaining model performance through knowledge distillation. Dong et al. (Dong et al., 2023) performed high-precision fitting for exponential and GeLU operations through piecewise polynomials. They successfully performed privacy-preserving inference on LLMs like LLaMA-7B. Although the existing SMPC technologies in LLMs still face challenges in efficiency, performance and cost, their prospects are extensive.
@@ 7.3.2. Security defense
Regarding adversarial example attacks and jailbreak attacks, existing countermeasures mainly consider output detection and processing, prompt engineering, and robustness training. We detail these defenses below.
@@
Output detection and processing. Some researchers detect and process malicious outputs during the generation phase to resist jailbreak attacks. Deng et al. (Deng et al., 2023) proved many closed-source LLMs have defense mechanisms, including keyword and semantic detection schemes. In addition, companies like Microsoft and NVIDIA have developed various detectors for harmful content. However, the training data limits classifier-based detection schemes, and adaptive jailbreak attacks can bypass them (Yang et al., 2024b). To improve detection effectiveness, OpenAI and Meta directly use GPT-4 and LLaMA2 to detect harmful content. Chen et al. (Chen et al., 2023) used multiple LLMs to resist jailbreak attacks. Specifically, they randomly selected responses from these LLMs as the output result.
@@
Prompt engineering. Users can manually or automatically optimize prompts, making LLMs better understand the context of these instructions. Therefore, some researchers used prompt engineering to eliminate the malicious goals of prompts, resulting in valuable and harmless responses. Li et al. (Li et al., 2023c) designed a purification scheme. They introduced random noise into the prompts and reconstructed them using a BERT-based mask language model. Robey et al. (Robey et al., 2023) found that jailbreak prompts are vulnerable to character-level perturbations. Therefore, they randomly perturbed multiple prompt copies and identified texts with high entropy as infected prompts. Jain et al. (Jain et al., 2023) proposed two prompt processing methods: paraphrasing and re-tokenization. Specifically, paraphrasing uses a generative model to modify the prompts, retaining the original instructions while removing adversarial tokens. Re-tokenization represents the prompts with multiple smaller tokens. It can disrupt the effect of malicious tokens while having minimal impact on clean samples. Wei et al. (Wei et al., 2023b) considered the few-shot scenario. They inserted a small number of defensive demonstrations into the prompts, mitigating jailbreak attacks and backdoor attacks.
@@
Robustness training. Developers can control the learning process to defend against various security attacks. Dong et al. (Dong et al., 2021) found that adversarial training can lead to catastrophic forgetting. They proposed a robust fine-tuning strategy to preserve the features learned by the LLM. Currently, most LLMs establish safety guardrails through the RLHF technology, defending against jailbreak attacks (Bai et al., 2022). Bianchi et al. (Bianchi et al., 2023) constructed a small amount of safety instructions (i.e., 3%) to improve the robustness of the LLaMA model. Sun et al. (Sun et al., 2024) argued that alignment tuning with human supervision was too costly. They leveraged another LLM to generate high-quality alignment instructions, constructing safety guardrails with minimal human supervision. Similarly, Shi et al. (Shi et al., 2023) generated high-quality preference data through three steps: reverse instruction tuning, instruction induction, and expert model evaluation, addressing the high labor costs in preference data annotation.
@@ 8. The Risks of Deploying LLM-based Agents
Some researchers leverage LLMs to construct agent systems, which are known as LLM-based agents. They are application systems that can understand natural language instructions and interact with humans. In general, LLM-based agents integrate various functional modules to perform complex tasks rather than merely running multiple rounds of queries. Currently, there are two application scenarios: single-agent systems and multi-agent systems. A multi-agent system consists of many LLM-based agents, each responsible for a specific task or role. For example, a health management system includes a data collection agent, an analysis agent, a report generation agent, and an interaction agent.
As illustrated in Figure 14, deploying LLM-based agents involves two malicious entities: users and agents. First, malicious users can manipulate prompts to launch backdoor and jailbreak attacks. Second, interactions between malicious agents pose privacy and security risks, such as unauthorized interactions and domino effects. Regarding these threats, we introduce privacy and security risks, respectively. Subsequently, we propose potential countermeasures to provide researchers with additional defenses, as shown in Figure 17.
@@ 8.1. Privacy risks of deploying LLM-based agents
Like the privacy risks discussed in Section 7.1, deploying LLM-based agents also faces prompt extraction attacks and various privacy attacks. We merely focus on privacy risks unique to LLM-based agents. Firstly, users inadvertently share their private information during interactions with LLM-based agents. The system may memorize user interaction history, allowing malicious users to steal private information through prompt engineering. To achieve this goal, malicious users can leverage jailbreak attacks and prompt injection attacks. Moreover, the multi-agent system faces more privacy risks due to high-frequency interactions. Different agents in the system undertake various roles and permissions. Some agents may access sensitive data beyond their permission scope or expose sensitive information to agents who should not access it. When performing collaborative tasks, multiple agents will share and process private information. An attacker can access private information across the system by compromising one agent, as shown in Figure 15.
Li et al. (Li et al., 2024b) found that data transmission between agents can be stolen, leading to privacy leakage. They also explored unauthorized interactions among agents, which may generate much private information. However, the interactions are not transparent, so it is hard to supervise this generated information, as illustrated in Figure 15. Therefore, it is essential to strictly control and transparently manage access to private data.
@@ 8.2. Security risks of deploying LLM-based agents
Yang et al. (Yang et al., 2024a) explored backdoor attacks for LLM-based agents, proposing query-attack, observation-attack and thought-attack. For the first manner, given a query with the trigger, the backdoored agents perform predefined malicious actions (e.g., send scam messages). For the second manner, adversaries can hide the trigger in interactions rather than input prompts. When the backdoored agent receives an observation result containing a trigger, it will perform malicious actions in subsequent steps. In the third manner, adversaries aim to alter the behaviors of specific steps while keeping benign actions. For instance, a backdoored agent might invoke a specific tool (e.g., Google Translate) under a trigger prompt. To achieve these attacks, adversaries need to create poisoned data and insert a backdoor into the LLM-based agent by training on it.
In addition to both attacks, malicious agents also pose security threats. For the single-agent system, attackers can modify the role settings of victim agents, causing them to exhibit harmful behaviors. Tian et al. (Tian et al., 2023) found that malicious agents can interact harmful content with other agents, affecting their behavior in a domino effect, as displayed in Figure 16. This risk improves the vulnerability of all LLM-based agents within a communication network. Meanwhile, the multi-agent system faces the threat of system-level attacks that can modify the role settings of the entire system.
@@ 8.3.1. Privacy protection
Potential defenses focus on memorized data and output results to address privacy leaks caused by malicious users. For memorized data, defenders can identify PII and use data masking techniques to hide sensitive information, thereby reducing the risk of PII leakage. For output results, defenders can implement filtering and auditing processes to prevent sensitive information from being transmitted to other entities. As introduced in Section 7.3.1, both rule-based and classifier-based detection schemes can be applied. Potential defenses include authority management and real-time feedback, which can mitigate privacy threats that agent interactions pose. In the first manner, defenders can establish clear controls for private data access, setting access permissions for different roles within multi-agent systems. For real-time feedback, defenders can dynamically monitor and adjust output results to mitigate privacy risks caused by unauthorized interactions.
While these countermeasures can effectively mitigate privacy risks discussed in Section 8.1, comprehensive research is lacking. Therefore, future efforts focus on targeted defenses against unique privacy threats for LLM-based agents.
@@ 8.3.2. Security defense
Agent processing. These countermeasures mainly address the security risks posed by malicious agents. Defenders can establish multi-level consistency frameworks in multi-agent systems, ensuring alignment with human values. For role attacks, defenders can design robust filters to detect malicious agents based on role attributes, improving system safety. In addition, Li et al. (Li et al., 2024b) found that a high-level agent guides its subordinate agents. Thus, constraining the high-level agent can prevent subordinate agents from producing harmful behaviors. These countermeasures can reduce the domino effect in multi-agent interactions.
While these countermeasures can effectively mitigate security risks discussed in Section 8.2, comprehensive research is also lacking. Comprehensive research is still needed to address evolving security threats for LLM-based agents.
@@ 9.1. Federated learning for LLMs
In practice, massive corpora come from many sources. Therefore, some researchers explore the distributed training framework. They assume that individual participants have local corpora and use federated learning to develop an LLM collaboratively. Federated learning can allow for the development of LLMs while protecting the privacy of all participants. For example, McMahan et al. (McMahan et al., 2017) improved the federated averaging algorithm, providing high-quality privacy protection for language models. However, pre-training LLMs consumes a lot of computing and communication resources. The federated learning applied to LLMs is used in fine-tuning. Zhang et al. (Zhang et al., 2023c) combined federated learning with Low-Rank Adaptation (LoRA), achieving efficient fine-tuning and privacy protection. Subsequently, Xu et al. (Xu et al., 2023) adopted differential privacy, partial embedding updates and LoRA to balance resource usage, privacy protection, and performance better. Similarly, Fan et al. (Fan et al., 2023) and Jiang et al. (Jiang et al., 2023b) focused on parameter-efficient federated fine-tuning methods for LLMs.
Although federated learning ensures that participants’ data do not leave the local environment, the updated gradient information during the collaboration process might make this framework susceptible to reconstruction and inference attacks. In this context, attackers (i.e., malicious participants) could access shared gradients of victim participants, inferring or reconstructing the private information. For reconstruction attacks, Balunovic et al. (Balunovic et al., 2022) assumed attackers could access the global model weights and the training data’s gradients. They modeled the prior probabilities using an auxiliary language model (e.g., GPT-3), and alternately applied continuous and discrete optimization algorithms to reconstruct the training data. Li et al. (Li et al., 2023b) reduced the strength of the assumption and designed a reconstruction attack that does not require the modification of the model parameters. They used gradient information and intermediate features to enhance the reconstruction effects. For inference attacks, some researchers found that such the risk could occur in federated learning framework (Melis et al., 2019). Specifically, attackers have some auxiliary data labeled with the target attribute. They could implement attribute inference attacks, either passively or actively, to infer whether victim participants’ data contains the attribute.